# Local Reinforcement Learning with Action-Conditioned Root Mean Squared Q-Functions

### v1

@@ 1 Introduction
The success of deep learning has relied on backpropagation (Rumelhart et al.,, 1986), a procedure that has significant limitations in terms of biological plausibility as it requires synchronous computations and weight symmetry. Many works have provided backprop-free alternatives for training deep neural networks (Lillicrap et al., 2016a, ; Nøkland,, 2016; Nøkland and Eidnes,, 2019). Notably, Hinton, (2022) proposed the Forward-Forward algorithm (FF), a new approach that performs layerwise contrastive learning between positive and negative samples. This algorithm is lightweight and entirely eliminates the need for backpropagation, thereby addressing some of the biological plausibility concerns.
However, most studies on backprop-free methods are focused on the search for a biologically plausible mechanism for performing gradient updates on supervised tasks. Could a biologically plausible source of learning signals be equally meaningful? Reward-centric environments and temporal-difference (TD) methods (Sutton,, 1988) serve as natural candidates for filling this gap. Biological brains have evolved through a series of reward-guided evolution, while ample evidence has shown that our brains could be implementing TD (Schultz et al., 1997a, ; O’Doherty et al.,, 2003; Watabe-Uchida et al.,, 2017; Amo et al.,, 2022). Since the goodness score in FF models the “compatibility” between the inputs and labels, this local learning paradigm can be readily adapted to a reinforcement learning (RL) setting where we model the value of an input state and an action from each layer’s activities. See Figure 1 for a comparison between the supervised learning and RL setups of the forward-forward learning paradigm.
Towards integrating local methods and RL, Guan et al., (2024) recently proposed Artificial Dopamine (AD) that incorporates top-down and temporal connections in a Q-learning framework. Since the local Q-function estimation needs to be explicitly predicted, Guan et al., (2024) uses a dot-product between two sets of mappings from the inputs that produces the value estimate for each action. This design, while backprop-free, makes the architecture more flexible modeling complex inputs. However, AD still relies on the output of the dot-product to be the same dimension as the action space, limiting the flexibility of the method.
Figure 1: Local learning paradigms inspired by the Forward-Forward (FF) algorithm (Hinton,, 2022). a) The original FF is designed for supervised learning, where each layer models the “goodness” between image $x$ and label $y$. Information is carried forward only through bottom-up and optionally top-down connections without backpropagation. b) We extend $F$ local learning for reinforcement learning—each layer takes a state observation $x$ and an action candidate $a$ as input, and estimates the Q value by taking the root mean squared function of the hidden vector.
@@
We evaluate our method on the MinAtar benchmark (Young and Tian,, 2019) and the DeepMind Control Suite, challenging suites designed to test RL algorithms in low-dimensional settings where local methods remain viable. Our results show that our method consistently outperforms current local RL methods and surpasses conventional backprop-based value-learning methods in most games, demonstrating strong decision-making capabilities without relying on backpropagation. Through this contribution, we seek to encourage further exploration of the intersection between RL and biologically plausible learning methods.
@@ 2 Related Works
In recent years, several backprop-free training algorithms have been proposed to address the limitations of traditional backpropagation in neural networks (Lillicrap et al., 2016a, ; Nøkland,, 2016; Nøkland and Eidnes,, 2019; Belilovsky et al.,, 2019; Baydin et al.,, 2022; Ren et al.,, 2023; Fournier et al.,, 2023; Singhal et al.,, 2023; Innocenti et al.,, 2025). One notable method is the Forward-Forward Algorithm (FF) (Hinton,, 2022), which offers a biologically plausible alternative to backpropagation. To extend the capabilities of FF, Ororbia and Mali, (2023) proposed the Predictive Forward-Forward Algorithm, showing that a top-down generative circuit can be trained jointly with FF. Tosato et al., (2023) found that models trained with FF objectives generate highly sparse representations. This pattern closely resembles the observations of neuronal ensembles in cortical sensory areas, suggesting FF may be a suitable candidate for modeling biological learning. Recently, Sun et al., (2025) proposed DeeperForward, integrating residual connections (He et al.,, 2016), the mean goodness function, and a channel-wise cross-entropy based objective function (Papachristodoulou et al.,, 2024) into FF. However, these works have mostly focused on supervised image classification rather than RL tasks.
@@
TD methods for value estimation have been particularly useful in the recent decade as the rise of deep neural networks offers a powerful function approximator. Mnih et al., (2015) introduced DQN, where a deep neural network is applied to approximate the Q-Function. They showed that this method significantly outperformed earlier methods on the Atari 2600 games, initiating a family of methods built upon this architecture (Van Hasselt et al.,, 2016; Wang et al.,, 2016; Dabney et al., 2018b, ; Hessel et al.,, 2018; Fortunato et al.,, 2018; Dabney et al., 2018a, ; Hausknecht and Stone,, 2015). In actor-critic architectures, it is also common to use a deep neural network for value and advantage estimation (Schulman et al.,, 2017; Schulman et al., 2015a, ; Schulman et al., 2015b, ; Lillicrap et al., 2016b, ; Mnih et al.,, 2016; Haarnoja et al., 2018b, ; Haarnoja et al., 2018a, ; Fujimoto et al.,, 2018; Gruslys et al.,, 2018; Abdolmaleki et al.,, 2018; Yarats et al., 2021b, ; Yarats et al., 2021a, ). For planning-based methods using either Monte Carlo tree search (MCTS) or a learned model, value estimation is also significant in driving the planning process (Schrittwieser et al.,, 2020; Silver et al.,, 2016, 2018; Hansen et al.,, 2024; Sacks et al.,, 2024; Ye et al.,, 2021; Hafner et al.,, 2025, 2021, 2020, 2019) Yet, few works have investigated the capability of local learning on value estimation.
@@
An important design choice in value estimation is whether the network is conditioned on the action. Early neural value estimation methods Riedmiller, (2005) incorporated action conditioning by incorporating both state and action as model inputs. With the advent of deep neural network approaches such as DQN, practices began to diverge. Purely value-based methods like DQN are typically only state-conditioned, with action-specific predictions produced at the output layer by indexing over action values. This design is computationally efficient and well-suited for discrete tasks with low-dimensional action spaces. In contrast, actor–critic methods developed for high-dimensional continuous control tasks Lillicrap et al., 2016b ; Haarnoja et al., 2018a condition on both state and action at the input of their critic networks. Although this distinction is largely arbitrary in backpropagation-based architectures and can be adapted to the task, we show that action conditioning at model inputs is strictly preferable for local RL.
@@
The concept of decentralized RL can be dated back to the dawn of RL. Klopf, (1982) introduced the idea of the hedonistic neuron, which hypothesized that each of our neurons may be guided by their independent rewards. Instead of being a miniscule part of a large operating neural system, each neuron may be an RL agent itself. In modern RL literature, the localized formulation of RL methods can be related to the multi-agent RL (MARL) setup, where multiple independent agents can be designed to cooperate well toward maximizing their joint rewards (Tan,, 1993; Foerster et al.,, 2017; Palmer et al.,, 2018; Su et al.,, 2022; Lauer and Riedmiller,, 2000; Jiang and Lu,, 2023; De Witt et al.,, 2020; Su and Lu,, 2022; Arslan and Yüksel,, 2016; Jin et al.,, 2022). Conveniently, we can frame the problem of training RL using local objectives as a MARL problem where each agent represents different modules within a network. Recently, Seyde et al., (2023) has explored a similar approach for the continuous control problem, showing that using a separate critics network for each fixed action after action discretization works surprisingly well. Guan et al., (2024) shows that a network with nonlinear local operations, decentralized objectives, and top-down connections across the temporal dimension can exceed state-of-the-art methods trained end-to-end. We extend upon this literature of decentralized methods for value estimation.
@@ 3 Background
The FF Algorithm (Hinton,, 2022), as its name denotes, uses two forward passes instead of one forward pass and one backward pass used in backpropagation. The first forward pass carries the positive data, or real data, while the second pass carries the negative data, or fake data either manually defined or synthetically generated by the network. The network is then trained by maximizing the goodness of each layer in the positive pass, while minimizing the goodness of each layer in the negative pass. The definition of goodness based on a hidden vector $z$ is as follows:
@@
Figure 2: High-level computation diagram between Guan et al., (2024) and ARQ. Key implementations of ARQ are highlighted in red. AD cells take activations (highlighted in blue, darker color means earlier layer) and the state observation as input and produces a vector of size $n_{a}$, each indicating the value prediction of an action candidate. Our ARQ takes activations, the state observation, and the action candidate as input, and produces a hidden vector of arbitrary size, before passing it through a root mean squared function to yield a scalar prediction.
A widely used class of methods for value estimation is temporal difference (TD) learning (Sutton,, 1988), which bootstraps value estimates by blending immediate rewards with future predictions, allowing for online, incremental updates. This method paved the way for the development of many subsequent approaches, particularly Q-learning. Take a q-function $Q(s,a)$. To update the function given an experience $(S_{t},a,r,S_{t+1)}$, Q-learning makes the following iterative update
@@
Recently, the rise of neural networks pushed q-learning to new heights. Mnih et al., (2015) proposed DQN, approximating q-function values using a deep neural network. Based on the Bellman equation, DQN constructs a mean squared error function as the objective, namely
@@
Mnih et al., (2015) tested their agents on the Atari 2600 environment, and show that a convolutional neural network trained in this fashion is able to achieve near-human performance level from raw pixel inputs, a feat previously considered far-fetched.
@@
AD (Guan et al.,, 2024) trains a local RL agent using Q-learning. An AD network is consisted of multiple AD cells, each of which makes an independent estimation of $Q(S_{t},A_{t})$. To yield a scalar estimation, each AD cell adopts an attention-like mechanism to compute a weighted sum of its hidden activations using weights from a separate linear projection, effectively incorporating nonlinearity while maintaining backprop-free. Additionally, each AD cell takes inputs from the layer below, the layer above, and also the raw state observation, enabling skip connections, top-down connections, and information flow throughout the temporal dimension in an RL environment. Mathematically, an AD cell at depth $l$ conducts the following operations,
@@ 4.1 ARQ
where $\theta$ denotes the parameters of the network , and $z$ denotes a hidden vector produced by the network.
To train this network, we update our weights using the same mean squared objective function as Deep Q-Learning (Mnih et al.,, 2015). Namely,
@@
ARQ can be implemented out-of-the-box in place of the standard Q-learning formulation. Given any intermediate vector produced by an arbitrary neural network architecture, ARQ can extract scalar statistics that serve as a prediction for the estimated Q-value without any parameters. This property allows architectures designed for local RL to enjoy greater flexibility.
@@
Due to the nature of goodness functions producing scalar values, it is natural to implement ARQ with action conditioning at the model input. Concretely, to estimate $Q_{\theta}(s,a)$, the neural network $\theta$ takes both the state vector $s$ and the action vector $a$ as inputs and outputs a single scalar prediction. This contrasts with implementations such as Mnih et al., (2015) and Guan et al., (2024), where the model receives only the state vector $s$ and produces an output of dimension $n_{a}$, with each entry corresponding to the value of a discrete action. We demonstrate in Section 5 that this minor design decision is critical to the performance of local RL methods. For tasks with discrete action spaces, we use a one-hot vector to represent an action candidate. For tasks with continuous action spaces, we apply bang-bang discretization on the action space following Seyde et al., (2021) and condition the network on the binary action vector.
@@ 4.2 Implementation
To evaluate our method against state-of-the-art local RL architectures, we implement AR on top of AD (Guan et al.,, 2024).
Our implementation is consisted of multiple cells stacked together, each of which takes inputs from the layer below, the layer above, the input observation, and an action candidate $a_{t}$ to make an estimation of $Q(s_{t},a_{t})$. Each cell adopts a similar attention-like mechanism as Guan et al., (2024). After the attention mechanism, we apply the goodness function on the intermediate vector after the attention computation. Specifically, a cell at depth $l$ conducts the following operations,
@@
$y_{t}^{l}=$ $\;\mathrm{tanh}(W_{att}X)^{T}h_{t}^{l},$ (13)
$\mu_{y}=\mathop{\mathbb{E}}_{y_{i}\in y}y_{i},\;\;$ $Q(s_{t},a_{t})=\sqrt{\mathop{\mathbb{E}}_{y_{i}\in y_{t}^{l}}(y_{i}-\mu_{y})^{2}},$ (14)
Gradients are passed only within each cell to ensure the architecture is backprop-free. Pseudocode comparising AD and ARQ can be found in Figure 3. Most training choices are inherited from Guan et al., (2024).
Algorithm 1 AD (Guan et al.,, 2024)
@@
3:Repeat $X$ along batch dim $n_{a}$ times
@@
Figure 3: Comparison of AD and ARQ implemented on top of AD. For ARQ, action conditioning is applied as part of the input (Line 4,5, Algorithm 2). Note that ARQ allows $Z_{2}$ and $y$ to have dimension $d$, which can be arbitrary, while AD fixes it at $n_{a}$, one for each action output.
@@
As demonstrated in Figure 3, ARQ allows the hidden output to have arbitrary dimensions. We hypothesize that ARQ’s flexibility to account for arbitrary hidden dimensions allows it to take full advantage of non-linearity within each AD cell. Furthermore, ARQ applies action conditioning at the model input, rather than using vector indices at the output layer as conditioning. We conjecture that this allows the entire module to produce representation specific to each state-action pair, rather than action-agnostic information based on only the observation. Combining these two properties, ARQ exploits the full capacity of the attention-like mechanism that modern local RL methods operates on, allowing greater expressivity of each state-action pair.

### v2

@@ 1 Introduction
The success of deep learning has relied on backpropagation (Rumelhart et al., 1986), a procedure that has significant limitations in terms of biological plausibility as it requires synchronous computations and weight symmetry. Many works have provided backprop-free alternatives for training deep neural networks (Lillicrap et al., 2016a; Nøkland, 2016; Nøkland and Eidnes, 2019). Notably, Hinton (2022) proposed the Forward-Forward algorithm (FF), a new approach that performs layerwise contrastive learning between positive and negative samples. This algorithm is lightweight and entirely eliminates the need for backpropagation, thereby addressing some of the biological plausibility concerns.
However, most studies on backprop-free methods are focused on the search for a biologically plausible mechanism for performing gradient updates on supervised tasks. Could a biologically plausible source of learning signals be equally meaningful? Reward-centric environments and temporal-difference (TD) methods (Sutton, 1988) serve as natural candidates for filling this gap. Biological brains have evolved through a series of reward-guided evolution, while ample evidence has shown that our brains could be implementing TD (Schultz et al., 1997a; O’Doherty et al., 2003; Watabe-Uchida et al., 2017; Amo et al., 2022). Since the goodness score in FF models the “compatibility” between the inputs and labels, this local learning paradigm can be readily adapted to a reinforcement learning (RL) setting where we model the value of an input state and an action from each layer’s activities. See Figure 1 for a comparison between the supervised learning and RL setups of the forward-forward learning paradigm.
Towards integrating local methods and RL, Guan et al. (2024) recently proposed Artificial Dopamine (AD) that incorporates top-down and temporal connections in a Q-learning framework. Since the local Q-function estimation needs to be explicitly predicted, Guan et al. (2024) uses a dot-product between two sets of mappings from the inputs that produces the value estimate for each action. This design, while backprop-free, makes the architecture more flexible modeling complex inputs. However, AD still relies on the output of the dot-product to be the same dimension as the action space, limiting the flexibility of the method.
Figure 1: Local learning paradigms inspired by the Forward-Forward (FF) algorithm (Hinton, 2022). a) The original FF is designed for supervised learning, where each layer models the “goodness” between image $x$ and label $y$. Information is carried forward only through bottom-up and optionally top-down connections without backpropagation. b) We extend $F$ local learning for reinforcement learning—each layer takes a state observation $x$ and an action candidate $a$ as input, and estimates the Q value by taking the root mean squared function of the hidden vector.
@@
We evaluate our method on the MinAtar benchmark (Young and Tian, 2019) and the DeepMind Control Suite, challenging suites designed to test RL algorithms in low-dimensional settings where local methods remain viable. Our results show that our method consistently outperforms current local RL methods and surpasses conventional backprop-based value-learning methods in most games, demonstrating strong decision-making capabilities without relying on backpropagation. Through this contribution, we seek to encourage further exploration of the intersection between RL and biologically plausible learning methods.
@@ 2 Related Works
In recent years, several backprop-free training algorithms have been proposed to address the limitations of traditional backpropagation in neural networks (Lillicrap et al., 2016a; Nøkland, 2016; Nøkland and Eidnes, 2019; Belilovsky et al., 2019; Baydin et al., 2022; Ren et al., 2023; Fournier et al., 2023; Singhal et al., 2023; Innocenti et al., 2025). One notable method is the Forward-Forward Algorithm (FF) (Hinton, 2022), which offers a biologically plausible alternative to backpropagation. To extend the capabilities of FF, Ororbia and Mali (2023) proposed the Predictive Forward-Forward Algorithm, showing that a top-down generative circuit can be trained jointly with FF. Tosato et al. (2023) found that models trained with FF objectives generate highly sparse representations. This pattern closely resembles the observations of neuronal ensembles in cortical sensory areas, suggesting FF may be a suitable candidate for modeling biological learning. Recently, Sun et al. (2025) proposed DeeperForward, integrating residual connections (He et al., 2016), the mean goodness function, and a channel-wise cross-entropy based objective function (Papachristodoulou et al., 2024) into FF. However, these works have mostly focused on supervised image classification rather than RL tasks.
@@
TD methods for value estimation have been particularly useful in the recent decade as the rise of deep neural networks offers a powerful function approximator. Mnih et al. (2015) introduced DQN, where a deep neural network is applied to approximate the Q-Function. They showed that this method significantly outperformed earlier methods on the Atari 2600 games, initiating a family of methods built upon this architecture (Van Hasselt et al., 2016; Wang et al., 2016; Dabney et al., 2018b; Hessel et al., 2018; Fortunato et al., 2018; Dabney et al., 2018a; Hausknecht and Stone, 2015). In actor-critic architectures, it is also common to use a deep neural network for value and advantage estimation (Schulman et al., 2017, 2015a, 2015b; Lillicrap et al., 2016b; Mnih et al., 2016; Haarnoja et al., 2018b, a; Fujimoto et al., 2018; Gruslys et al., 2018; Abdolmaleki et al., 2018; Yarats et al., 2021b, a). For planning-based methods using either Monte Carlo tree search (MCTS) or a learned model, value estimation is also significant in driving the planning process (Schrittwieser et al., 2020; Silver et al., 2016, 2018; Hansen et al., 2024; Sacks et al., 2024; Ye et al., 2021; Hafner et al., 2025, 2021, 2020, 2019) Yet, few works have investigated the capability of local learning on value estimation.
@@
An important design choice in value estimation is whether the network is conditioned on the action. Early neural value estimation methods Riedmiller (2005) incorporated action conditioning by incorporating both state and action as model inputs. With the advent of deep neural network approaches such as DQN, practices began to diverge. Purely value-based methods like DQN are typically only state-conditioned, with action-specific predictions produced at the output layer by indexing over action values. This design is computationally efficient and well-suited for discrete tasks with low-dimensional action spaces. In contrast, actor–critic methods developed for high-dimensional continuous control tasks Lillicrap et al. (2016b); Haarnoja et al. (2018a) condition on both state and action at the input of their critic networks. Although this distinction is largely arbitrary in backpropagation-based architectures and can be adapted to the task, we show that action conditioning at model inputs is strictly preferable for local RL.
@@
The concept of decentralized RL can be dated back to the dawn of RL. Klopf (1982) introduced the idea of the hedonistic neuron, which hypothesized that each of our neurons may be guided by their independent rewards. Instead of being a miniscule part of a large operating neural system, each neuron may be an RL agent itself. In modern RL literature, the localized formulation of RL methods can be related to the multi-agent RL (MARL) setup, where multiple independent agents can be designed to cooperate well toward maximizing their joint rewards (Tan, 1993; Foerster et al., 2017; Palmer et al., 2018; Su et al., 2022; Lauer and Riedmiller, 2000; Jiang and Lu, 2023; De Witt et al., 2020; Su and Lu, 2022; Arslan and Yüksel, 2016; Jin et al., 2022). Conveniently, we can frame the problem of training RL using local objectives as a MARL problem where each agent represents different modules within a network. Recently, Seyde et al. (2023) has explored a similar approach for the continuous control problem, showing that using a separate critics network for each fixed action after action discretization works surprisingly well. Guan et al. (2024) shows that a network with nonlinear local operations, decentralized objectives, and top-down connections across the temporal dimension can exceed state-of-the-art methods trained end-to-end. We extend upon this literature of decentralized methods for value estimation.
@@ 3 Background
The FF Algorithm (Hinton, 2022), as its name denotes, uses two forward passes instead of one forward pass and one backward pass used in backpropagation. The first forward pass carries the positive data, or real data, while the second pass carries the negative data, or fake data either manually defined or synthetically generated by the network. The network is then trained by maximizing the goodness of each layer in the positive pass, while minimizing the goodness of each layer in the negative pass. The definition of goodness based on a hidden vector $z$ is as follows:
@@
Figure 2: High-level computation diagram between Guan et al. (2024) and ARQ. Key implementations of ARQ are highlighted in red. AD cells take activations (highlighted in blue, darker color means earlier layer) and the state observation as input and produces a vector of size $n_{a}$, each indicating the value prediction of an action candidate. Our ARQ takes activations, the state observation, and the action candidate as input, and produces a hidden vector of arbitrary size, before passing it through a root mean squared function to yield a scalar prediction.
A widely used class of methods for value estimation is temporal difference (TD) learning (Sutton, 1988), which bootstraps value estimates by blending immediate rewards with future predictions, allowing for online, incremental updates. This method paved the way for the development of many subsequent approaches, particularly Q-learning. Take a q-function $Q(s,a)$. To update the function given an experience $(S_{t},a,r,S_{t+1)}$, Q-learning makes the following iterative update
@@
Recently, the rise of neural networks pushed q-learning to new heights. Mnih et al. (2015) proposed DQN, approximating q-function values using a deep neural network. Based on the Bellman equation, DQN constructs a mean squared error function as the objective, namely
@@
Mnih et al. (2015) tested their agents on the Atari 2600 environment, and show that a convolutional neural network trained in this fashion is able to achieve near-human performance level from raw pixel inputs, a feat previously considered far-fetched.
@@
AD (Guan et al., 2024) trains a local RL agent using Q-learning. An AD network is consisted of multiple AD cells, each of which makes an independent estimation of $Q(S_{t},A_{t})$. To yield a scalar estimation, each AD cell adopts an attention-like mechanism to compute a weighted sum of its hidden activations using weights from a separate linear projection, effectively incorporating nonlinearity while maintaining backprop-free. Additionally, each AD cell takes inputs from the layer below, the layer above, and also the raw state observation, enabling skip connections, top-down connections, and information flow throughout the temporal dimension in an RL environment. Mathematically, an AD cell at depth $l$ conducts the following operations,
@@ 4.1 ARQ
where $\theta$ denotes the parameters of the network and $z$ denotes a hidden vector produced by the network.
To train this network, we update our weights using the same mean squared objective function as previous Q-learning methods (Mnih et al., 2015). Namely,
@@
ARQ can be implemented out-of-the-box in place of the standard Q-learning formulation. Given any intermediate vector produced by an arbitrary neural network architecture, ARQ can extract scalar statistics that serve as a prediction for the estimated value without any parameters. This property allows architectures designed for local RL to enjoy greater flexibility.
@@
Due to the nature of goodness functions to produce scalar values, it is natural to implement ARQ with action conditioning at the model input. Concretely, to estimate $Q_{\theta}(s,a)$, the neural network $\theta$ takes both the state vector $s$ and the action vector $a$ as inputs and outputs a single scalar prediction. This contrasts with implementations such as Mnih et al. (2015) and Guan et al. (2024), where the model receives only the state vector $s$ and produces an output of dimension $n_{a}$, with each entry corresponding to the value of a discrete action. We demonstrate in Section 5 that this minor design decision is critical to the performance of local RL methods. For tasks with discrete action spaces, we use a one-hot vector to represent an action candidate. For tasks with continuous action spaces, we apply bang-bang discretization on the action space following Seyde et al. (2021) and condition the network on the binary action vector.
@@ 4.2 Implementation
To evaluate our method against state-of-the-art local RL architectures, we implement AR on top of Guan et al. (2024).
Our implementation is consisted of multiple cells stacked together, each of which takes inputs from the layer below, the layer above, the input observation, and an action candidate $a_{t}$ to make an estimation of $Q(s_{t},a_{t})$. Each cell adopts a similar attention-like mechanism as Guan et al. (2024). After the attention mechanism, we apply the goodness function on the intermediate vector after the attention computation. Specifically, a cell at depth $l$ conducts the following operations,
@@
$y_{t}^{l}=$ $\;\mathrm{tanh}(X^{T}W_{att_{2}}^{T}W_{att_{1}}X)h_{t}^{l},$ (13)
$\mu_{y}=\mathop{\mathbb{E}}_{y_{i}\in y_{t}^{l}}y_{i},$ $\;\;Q(s_{t},a_{t})=\sqrt{\mathop{\mathbb{E}}_{y_{i}\in y_{t}^{l}}(y_{i}-\mu_{y})^{2}},$ (14)
Gradients are passed only within each cell to ensure the architecture is backprop-free.
Pseudocode for AD and ARQ is provided in Figure 3. In both architectures, the intermediate quantities $Z_{1}$, $Z_{2}$, and $h_{t}^{l}$ play roles analogous to the query, key, and value vectors in self-attention. The computation of $Z_{2}^{T}Z_{1}$ (Line 8, Algorithm 2) produces a dimension-wise interaction map that determines how information is redistributed across latent dimensions, similar in spirit to an attention mechanism but applied over feature dimensions rather than token positions. The key distinction in ARQ is that $Z_{2}$ is not restricted to have width $n_{a}$; instead, its dimensionality can be chosen freely. This flexibility allows ARQ to learn richer state–action interactions than AD, whose dimensionality is constrained by the cardinality of the action space.
Algorithm 1 AD (Guan et al., 2024)
@@
3:   $\triangleright$ Dimension: $d$
4:Repeat $X$ along batch dim $n_{a}$ times
@@
Figure 3: Comparison of AD and ARQ implemented on top of AD. For ARQ, action conditioning is applied as part of the input (Line 5,6, Algorithm 2). Note that ARQ allows $Z_{2}$ and $y$ to have dimension $d$ (red), which can be arbitrary, while AD fixes it at $n_{a}$ (blue), one for each action output.
@@
As demonstrated in Figure 3, ARQ allows the hidden output to have arbitrary dimensions. We conjecture that ARQ’s flexibility to account for arbitrary hidden dimensions allows it to take full advantage of non-linearity within each AD cell. Furthermore, ARQ applies action conditioning at the model input, rather than using vector indices at the output layer as conditioning. We conjecture that this allows the entire module to produce representation specific to each state-action pair, rather than action-agnostic information based on only the observation. Combining these two properties, ARQ exploits the full capacity of the attention-like mechanism that modern local RL methods operates on, allowing greater expressivity of each state-action pair.